Uberville Drivers Inc. reported $750,000 in sales in the recent year, $650,000 in costs and $10,000 in interest. The frms tax rate is 34%, and recent years dividend payment was $34,400. The company is planning to grow by 20% in the next year. Costs are proportional to sales. Interest is not. The company will face the same tax rate, and is targeting the same dividend payout ratio Construct a pro-forma income statement. What is the pro-forma addition to retained earnings? The firm has $538,000 in assets and $307,000 in equity. Accounts payable are $65,000, notes payable are $16,000 and long term debt is $150,000. Assets and accounts payable are proportional to sales. All other balance sheet items are not. Construct a pro-forma balance sheet. What are the External Funds Needed? the current debt-to-equity ratio after its growth. What part of the external funds needed Suppose that the firm would like to maintain will it raise in equity?
Answer #1

1.

Uberville Drivers Inc.
Proforma Income Statement
Sales 900,000
Less: Costs 780,000
Earnings before Interest and Taxes 120,000
Less: Interest Expense 10,000
Earnings before Taxes 110,000
Taxes ( 34 % ) 37,400
Net Income 72,600
Less: Dividends 42,044
Addition to Retained Earnings 30,556

2. Total Assets : $ 538,000 x 1.20 = $ 645,600

Equity : $ 307,000 + $ 30,556 = $ 337,556

Accounts Payable: $ 65,000 x 1.20 = $ 78,000.

Uberville Drivers Inc.
Proforma Balance Sheet
Assets $ 645,600 Accounts Payable $ 78,000
Notes Payable 16,000
Long-term Liabilities 150,000
Equity 337,556
Total Assets $ 645,600 Total Liabilities and Equity

$ 581,556

External Funds Needed : $ 645,600 - $ 581,556 = $ 64,044.

3. Current Debt: Equity Ratio + $ ( 65,000 + 16,000 + 150,000) / $ 307,000 = 0.7524 : 1

Let the value of equity be y.

0.7524 y + y = 64,044

y = 64,044 / 1.7524 = $ 36,546.45

Amount to be raised in equity : $ 36,546.